The cost in millions to remove x% of the pollutant in an ocean modeled by C=6000/(200-2x):
What is the cost to remove 75% of the pollutant?
What is the cost to remove 90% of the pollutant?
What is the cost to remove 99% of the pollutant?

Answer by checkley77(12569)

C=6000/(200-2x)
C=6000/(200-2*75)
C=6000/(200-150)
C=6000/50
C=$120 MILLION.
C=6000/(200-2*90)
C=6000/(200-180)
C=6000/20
C=$300 MILLION.
C=6000/(200-2*99)
C=6000/200-198)
C=6000/2
C=$3,000 MILLION
What is the cost to remove 75% of the pollutant? $120 MILLION.
What is the cost to remove 90% of the pollutant? $300 MILLION.
What is the cost to remove 99% of the pollutant? $3,000 MILLION.